# Affordance-Aware Object Insertion via Mask-Aware Dual Diffusion

### v1

@@ Abstract
As a common image editing operation, image composition involves integrating foreground objects into background scenes. In this paper, we expand the application of the concept of Affordance from human-centered image composition tasks to a more general object-scene composition framework, addressing the complex interplay between foreground objects and background scenes. Following the principle of Affordance, we define the affordance-aware object insertion task, which aims to seamlessly insert any object into any scene with various position prompts. To address the limited data issue and incorporate this task, we constructed the SAM-FB dataset, which contains over 3 million examples across more than 3,000 object categories. Furthermore, we propose the Mask-Aware Dual Diffusion (MADD) model, which utilizes a dual-stream architecture to simultaneously denoise the RGB image and the insertion mask. By explicitly modeling the insertion mask in the diffusion process, MADD effectively facilitates the notion of affordance. Extensive experimental results show that our method outperforms the state-of-the-art methods and exhibits strong generalization performance on in-the-wild images. Please refer to our code on https://github.com/KaKituken/affordance-aware-any.
@@ 1 Introduction
In the image composition scenario, common sense guides our perception of the authenticity of synthesized images: a person cannot levitate, a water bottle needs a surface for support, and a boat should be floating in the water rather than on the ground. Deviation from such common sense often leads to semantic inconsistencies in synthesized images. From a composition perspective, the background’s semantic richness plays a pivotal role in defining the placement and characteristics of foreground objects. To better describe the influence of background semantics on the foreground, we borrow the concept of affordance into object-scene composition tasks. Previously, Kulal et al. [24] explored the concept of human affordance for image synthesis. Their work, focused on inserting humans into masked scenes, extends beyond mere color or view adjustments. To generalize the setting to arbitrary object-scene synthesis, we term the new task Affordance-Aware Object Insertion. This task challenges models to identify suitable locations and make necessary adjustments to foreground objects, ensuring the generated images adhere to physical laws.
@@
This work aims to build a foundation model for affordance-aware object insertion, which can put any object into any scene as shown in Figure LABEL:fig:teaser. There are three primary challenges involved. First, the model must accurately recognize the appropriate affordance relationship between a background and the foreground object to be inserted. Adjustments to the inserted foreground object are crucial for achieving the intended semantic consistency. The second challenge lies in the model’s ability to generalize across a diverse range of foreground objects. Previous generative image editing methods, such as textual inversion [11], DreamBooth [40], and DreamEdit [26], are subject-specific thus hard to generalize. Our goal is to train a model that can generalize to any object. Third, our model should support a variety of input prompts for users to specify insertion locations, ranging from sparse formats like points and bounding boxes to dense masks. Moreover, in the absence of explicit prompts, the model can autonomously determine appropriate insertion locations by analyzing the semantic content of both the background and foreground.
@@
We address these challenges through three components: task, data, and model. Extending the concept of affordance beyond Kulal et al.’s initial scope [24], the task of affordance-aware object insertion aims to place an arbitrary object into any scene, accommodating different positional prompts, even in the absence of explicit positional cues. It has significant implications for applications such as automated dataset synthesis. To support this task, a large-scale dataset is necessary. Existing image composition datasets, such as DreamEditBench [26], are limited in terms of the diversity of foreground object categories and the number of training samples. To overcome these limitations, we curate a new dataset called SAM-FB which is derived from SA-1B [23] for affordance learning. SAM-FB contains a variety of foreground object categories and over 3 million samples. With SAM-FB, we further propose a Mask-Aware Dual Diffusion (MADD) model to utilize the large-scale data, which is a diffusion-based framework that facilitates the seamless integration of diverse objects into any scene. During the denoising procedure, object position is progressively refined while the target RGB image is synthesized simultaneously, ensuring accurate alignment between objects and positions to achieve affordance-aware insertion. Furthermore, we present a unified representation for sparse and dense prompts, enabling our model to interpret and respond to various types of position inputs effectively.
@@
  • We formulate the task of affordance-aware object insertion, which extends general object-scene composition with affordance guidance, aiming to achieve realistic object insertion with diverse prompts.
  • We introduce SAM-FB, a large-scale dataset with diverse object categories for affordance-aware object insertion, which consists of over 3 million samples.
  • To facilitate the learning of the affordance concept, we propose the mask-aware dual diffusion model, which adopts a dual-stream architecture to simultaneously denoise the appearance and the insertion mask.

### v2

@@ Abstract
As a common image editing operation, image composition involves integrating foreground objects into background scenes. In this paper, we expand the application of the concept of Affordance from human-centered image composition tasks to a more general object-scene composition framework, addressing the complex interplay between foreground objects and background scenes. Following the principle of Affordance, we define the affordance-aware object insertion task, which aims to seamlessly insert any object into any scene with various position prompts. To address the limited data issue and incorporate this task, we constructed the SAM-FB dataset, which contains over 3 million examples across more than 3,000 object categories. Furthermore, we introduce a strong baseline: a Mask-Aware Dual Diffusion (MADD) model, which utilizes a dual-stream architecture to simultaneously denoise the RGB image and the insertion mask. By explicitly modeling the insertion mask in the diffusion process, MADD effectively facilitates the notion of affordance. Experimental results show that our method outperforms the state-of-the-art methods and exhibits strong generalization performance on in-the-wild images. Please refer to our code on https://github.com/KaKituken/affordance-aware-any.
@@ 1 Introduction
In the image composition scenario, common sense guides our perception of the authenticity of synthesized images: a person cannot levitate, a water bottle needs a surface for support, and a boat should be floating in the water rather than on the ground. Deviation from such common sense often leads to semantic inconsistencies on synthesized images. From a composition perspective, the background’s semantic richness plays a pivotal role in defining the placement and characteristics of foreground objects. To better describe the influence of background semantics on the foreground, we borrow the concept of affordance into object-scene composition tasks. Previously, Kulal et al. [24] explored the concept of human affordance for image synthesis. Their work, focused on inserting humans into masked scenes, extends beyond mere color or view adjustments. To generalize the setting to arbitrary object-scene synthesis, we term the new task Affordance-Aware Object Insertion. This task challenges models to identify suitable locations and make necessary adjustments to foreground objects, ensuring the generated images adhere to physical laws.
@@
This work aims to build a foundation model for affordance-aware object insertion, which can put any object into any scene as shown in Figure LABEL:fig:teaser. There are three primary challenges involved. First, the model must accurately recognize the appropriate affordance relationship between a background and the foreground object to be inserted. Adjustments to the inserted foreground object is crucial for achieving the intended semantic consistency. The second challenge lies in the model’s ability to generalize across a diverse range of foreground objects. Previous generative image editing methods, such as textual inversion [12], DreamBooth [43], and DreamEdit [27], are subject-specific thus hard to generalize. Our goal is to train a model that can generalize to any object. Third, our model should support a variety of input prompts for users to specify insertion locations, ranging from sparse formats like points and bounding boxes to dense masks. Moreover, in the absence of explicit prompts, the model can autonomously determine appropriate insertion locations by analyzing the semantic content of both the background and foreground.
@@
We address these challenges through three components: task, data, and model. Extending the concept of affordance beyond Kulal et al.’s initial scope [24], the task of affordance-aware object insertion aims to place an arbitrary object into any scene, accommodating different positional prompts, even in the absence of explicit positional cues. It has significant implications for applications such as automated dataset synthesis. To support this task, a large-scale dataset is necessary. Existing image composition datasets, such as DreamEditBench [27], are limited in terms of the diversity of foreground object categories and the number of training samples. To overcome these limitations, we curate a new dataset called SAM-FB which is derived from SA-1B [23] for affordance learning. SAM-FB contains a variety of foreground object categories and over 3 million samples. With SAM-FB, we further introduce a strong baseline: a Mask-Aware Dual Diffusion (MADD) model to utilize the large-scale data, which is a diffusion-based framework that facilitates the seamless integration of diverse objects into any scene. During the denoising procedure, object position is progressively refined while the target RGB image is synthesized simultaneously, ensuring accurate alignment between objects and positions to achieve affordance-aware insertion. Furthermore, we present a unified representation for sparse and dense prompts, enabling our model to effectively respond to various types of position inputs.
@@
  • We introduce affordance-aware object insertion, extending object-scene composition with affordance guidance to enable realistic insertions across diverse prompts.
  • We present SAM-FB, a large-scale dataset with over 3 million samples spanning diverse object categories to support affordance-aware insertion.
  • We propose a strong baseline model with a dual-stream architecture that denoises object appearance and the insertion mask, facilitating affordance learning.
@@ 2 Related Work
Model Task Dataset Sample Category Affordance Consistency
 [27] Customized Image Composition DreamEditBench 440 22 High (Real Paired)
 [34] Customized Image Composition MureCom 640 32 High (Real Paired)
 [49],  [44] Image-guidance Object Insertion Open-Images-v4 1.9M 600 Low (Masked)
 [30] Image-guidance Object Insertion COCO2014 165K 80 Low (Masked)
 [24] Human Affordance Insertion Private 2.4M 1 High (Video)
MADD (ours) Object Affordance Insertion SAM-FB 3.2M 3,439 High (Inpainted)
Table 1: Dataset choice and comparison. We inspect various potential datasets for affordance insertion. SAM-FB contains significantly more samples and object categories.    DreamEdit,    DreamCom,    PBE,    ObjectStitch,   GLIGEN,    Human Affordance.